1. Which of the following statements is least likely to be correct?

A. Securitizers are brokers that help to execute large orders.

B. Depository institutions include banks, credit unions, and savings and loans.

C. Alternative trading systems that do not reveal current client orders are known to be dark pools.

2. At the start of the year, an investor purchased 1000 shares of HSBC at the market price of $80 on full margin. The leverage ratio is 2.5 and the call money rate is 3%. Commission is charged at 0.25% on total trade value. What is the return to the investor if he sold the HSBC shares when the price is $85 at year end?

A. 9.77%

B. 9.84%

C. 15.6%

3. An investor purchased 1000 shares of HSBC at the market price of $80 on full margin. The initial margin requirement is 40% and the maintenance margin requirement is 30%. At what share price will an investor get a margin call?

A. $24.00

B. $68.57

C. $86.15
Answer #1

1> B is correct. Depository institutions accept depots, extend loans & provide withdrawals. Thus banks, credit unions and savings institutions are depository institutions

2> B . Return computation is closest to 9.84%

3>B. Account balance as per initial margin =1000* (80* 0.4) = 32000

maintanence margin = 1000* (80*0.3) = 24000

at share price $68.57 value of holding will be 80-68.57*1000=11430

Thus fall in value of 11430 (32000-11430) will make margin balance go lower than required .i.e. maintainence level (24000)

Thus at share price (B) margin call will be triggered